# TAPAAL HyperLTL: A Tool for Checking Hyperproperties of Petri Nets

### v1

@@ 2.1 Petri nets
Figure 1(b) shows an example of a P/T net where places from the set $P=\{v_{0},\ldots,v_{3},a(t_{0}),a(t^{\prime}_{0}),\ldots,a(t_{4}),a(t^{\prime}_{4})\}$ are drawn as circles, transitions from the set $T=\{t_{1},t^{\prime}_{1},\ldots,t_{4},t^{\prime}_{4},v_{1}^{\mathit{deliver}},v_{2}^{\mathit{deliver}}\}$ are drawn as rectangles, and arcs are the directed edges connecting either places to transitions or transitions to places. Unless otherwise stated, the default weight of all arcs is 11.
@@ 2.2 HyperLTL
Coming back to our example in Figure 1, the formula $\varphi_{1}$ (resp. $\varphi_{3}$) from Figure 1(c) expresses that there are two traces that both eventually (but possibly at different positions) receive a token in $v_{1}^{\prime}$ (resp. $v_{2}^{\prime}$) and each transition $t$ is fired in at most one of the traces (hence there must be a token in $a(t)$ in at least one of the two traces). In other words, $\varphi_{1}$ and $\varphi_{3}$ express that there are two disjoint paths in the graph in Figure 1(a), starting at $v_{0}$ and leading to $v_{1}$ resp. $v_{2}$. Analogously, $\varphi_{2}$ and $\varphi_{4}$ express similar properties, but requiring the existence of three disjoint paths. Hence, $\varphi_{1}$, $\varphi_{2}$, and $\varphi_{3}$ are satisfied by the example net NN, but not $\varphi_{4}$.
@@ 4 Congestion Case Study
In our concrete example, the self-composition does not require such a complicated synchronization mechanism as for each quantified trace we can create a copy of the net that can run completely concurrently (avoding the lock-step synchronization), while checking for the congestion using the shared places $a(t)$ that contain as many tokens as is the edge capacity. Figure 3 shows our simplified self-composition for our running example as well as a classical LTL query that expresses the same property as the HyperLTL formula $\varphi_{1}$ from Figure 1(c). To verify the classical LTL formula on the self-composed system, we benchmark our tool against the LTL engine of TAPAAL [31], the winner in the 2025 Model Checking Contest [33] in the LTL category. A ratio plot is depicted in Figure 4 where the $x$-axis contains all queries solved by at least one of the methods, sorted by the ratio of self-composition running time divided by TAPAAL HyperLTL running time. We remark that because the plot contains three different problem instances where the tools solve different number of queries (the numbers of answers in the parenthesis show the total number of solved problems by at least one of the methods), we use the percentage scale on the x-axis instead of the absolute count. For two disjoint paths, both methods are comparable, however, for 33 and 44 paths there is a clear advantage of using our new HyperLTL implementation. For example, for 44 paths, the self-composition timeouts (depicted by the straight horizontal line) on more than 50% of all queries that the HyperLTL implementation managed to solve. This is in particular true for queries with positive answers, as the on-the-fly method that we implemented in TAPAAL HyperLTL is more efficient than self-composition, where the net size explodes with number of trace variables in the HyperLTL formula.
@@ 5.1 TAPAAL HyperLTL Encoding
To encode the latency problem in TAPAAL HyperLTL, we assume as before a natural way to model packet hops using transition firings. The Petri net model for our running example from Figure 1(a) where the source node is $v_{0}$ and the target node is $v_{2}$ is presented in Figure 7. Additionally, in order to guarantee loop-freedom, for every place $p$ we add a new place $once(p)$ with a single token in its initial marking, such that any hop (transition firing) that places a token to $p$ will also consume a token from $once(p)$. This guarantees that we never visit the same place twice. To this end, all transitions that add tokens to the source place $v_{0}$ or remove tokens from the target place $v_{2}$ are also removed. In order to measure the latency, we add a new place $\mathit{counter}$ where every packet hop deposits a number of tokens that corresponds to the latency of the hop. For example firing $t_{2}$ deposits 33 tokens to the place $\mathit{counter}$ to represent latency 33 of the hop between the nodes $v_{0}$ and $v_{3}$. The HyperLTL formula checking if there are two loop-free traces that reach $v_{2}$ with a latency difference larger than or equal to $\ell$ is given under the net in Figure 7.
@@ 5.2 MCHyper Encoding
The basic network topology is encoded in the same way as in the previous section. We use two different encodings of the accumulated latency counter: the counter values are encoded either in binary or in unary. In the binary encoding we use for efficiency reasons three counters $C_{1}$, $C_{2}$ and $C_{3}$. The first counter $C_{1}$ uses $\lceil\text{log}_{2}(\ell)\rceil$ bits and stores the remainder of the accumulated latency divided by the latency bound $\ell$. The second counter $C_{2}$ uses $\lceil\text{log}_{2}(\frac{w_{\text{max}}}{\ell}|E|)\rceil$ bits where $w_{\text{max}}$ is the maximum latency on a link and stores the quotient of the division. For technical reasons, we also have a counter $C_{3}$; its value is always equal to the value of $C_{2}$ plus 11.
@@
$\exists\pi_{1}.\ \exists\pi_{2}.\\ \ \pi_{1}.goal\land\pi_{2}.goal\land((\pi_{1}.C_{2}>\pi_{2}.C_{3})\lor(\pi_{1}.C_{2}=\pi_{2}.C_{3}\land\pi_{1}.C_{1}\geq\pi_{2}.C_{1}))\ .$
@@ 5.3 Experimental Setup
Figure 9 shows the CPU time (on y-axis) for all independently sorted instances that where solved by TAPAAL HyperLTL and MCHyper. For each tool, we consider three benchmarks: one with $\ell=8$ and where all latency constants are between 1 to 5 and two more where all constants as well as $\ell$ are multiplied by 1515 respectively 255255. This multiplication creates instances that are equivalent to the basic one, however, the numbers in the model become larger.

### v2

@@ 2.1 Petri nets
Figure 1(b) shows an example of a P/T net where places from the set $P=\{v_{0},\ldots,v_{3},a(t_{0}),a(t^{\prime}_{0}),\ldots,a(t_{4}),a(t^{\prime}_{4})\}$ are drawn as circles, transitions from the set $T=\{t_{1},t^{\prime}_{1},\ldots,t_{4},t^{\prime}_{4},v_{1}^{\mathit{deliver}},v_{2}^{\mathit{deliver}}\}$ are drawn as rectangles, and arcs are the directed edges connecting either places to transitions or transitions to places. Unless otherwise stated, the default weight of all arcs is $1$.
@@ 2.2 HyperLTL
Coming back to our example in Figure 1, the formula $\varphi_{1}$ (resp. $\varphi_{3}$) from Figure 1(c) expresses that there are two traces that both eventually (but possibly at different positions) receive a token in $v_{1}^{\prime}$ (resp. $v_{2}^{\prime}$) and each transition $t$ is fired in at most one of the traces (hence there must be a token in $a(t)$ in at least one of the two traces). In other words, $\varphi_{1}$ and $\varphi_{3}$ express that there are two disjoint paths in the graph in Figure 1(a), starting at $v_{0}$ and leading to $v_{1}$ resp. $v_{2}$. Analogously, $\varphi_{2}$ and $\varphi_{4}$ express similar properties, but requiring the existence of three disjoint paths. Hence, $\varphi_{1}$, $\varphi_{2}$, and $\varphi_{3}$ are satisfied by the example net $N$, but not $\varphi_{4}$.
@@ 4 Congestion Case Study
In our concrete example, the self-composition does not require such a complicated synchronization mechanism as for each quantified trace we can create a copy of the net that can run completely concurrently (avoding the lock-step synchronization), while checking for the congestion using the shared places $a(t)$ that contain as many tokens as is the edge capacity. Figure 3 shows our simplified self-composition for our running example as well as a classical LTL query that expresses the same property as the HyperLTL formula $\varphi_{1}$ from Figure 1(c). To verify the classical LTL formula on the self-composed system, we benchmark our tool against the LTL engine of TAPAAL [31], the winner in the 2025 Model Checking Contest [33] in the LTL category. A ratio plot is depicted in Figure 4 where the $x$-axis contains all queries solved by at least one of the methods, sorted by the ratio of self-composition running time divided by TAPAAL HyperLTL running time. We remark that because the plot contains three different problem instances where the tools solve different number of queries (the numbers of answers in the parenthesis show the total number of solved problems by at least one of the methods), we use the percentage scale on the x-axis instead of the absolute count. For two disjoint paths, both methods are comparable, however, for $3$ and $4$ paths there is a clear advantage of using our new HyperLTL implementation. For example, for $4$ paths, the self-composition timeouts (depicted by the straight horizontal line) on more than 50% of all queries that the HyperLTL implementation managed to solve. This is in particular true for queries with positive answers, as the on-the-fly method that we implemented in TAPAAL HyperLTL is more efficient than self-composition, where the net size explodes with number of trace variables in the HyperLTL formula.
@@ 5.1 TAPAAL HyperLTL Encoding
To encode the latency problem in TAPAAL HyperLTL, we assume as before a natural way to model packet hops using transition firings. The Petri net model for our running example from Figure 1(a) where the source node is $v_{0}$ and the target node is $v_{2}$ is presented in Figure 7. Additionally, in order to guarantee loop-freedom, for every place $p$ we add a new place $once(p)$ with a single token in its initial marking, such that any hop (transition firing) that places a token to $p$ will also consume a token from $once(p)$. This guarantees that we never visit the same place twice. To this end, all transitions that add tokens to the source place $v_{0}$ or remove tokens from the target place $v_{2}$ are also removed. In order to measure the latency, we add a new place $\mathit{counter}$ where every packet hop deposits a number of tokens that corresponds to the latency of the hop. For example firing $t_{2}$ deposits $3$ tokens to the place $\mathit{counter}$ to represent latency $3$ of the hop between the nodes $v_{0}$ and $v_{3}$. The HyperLTL formula checking if there are two loop-free traces that reach $v_{2}$ with a latency difference larger than or equal to $\ell$ is given under the net in Figure 7.
@@ 5.2 MCHyper Encoding
The basic network topology is encoded in the same way as in the previous section. We use two different encodings of the accumulated latency counter: the counter values are encoded either in binary or in unary. In the binary encoding we use for efficiency reasons three counters $C_{1}$, $C_{2}$ and $C_{3}$. The first counter $C_{1}$ uses $\lceil\text{log}_{2}(\ell)\rceil$ bits and stores the remainder of the accumulated latency divided by the latency bound $\ell$. The second counter $C_{2}$ uses $\lceil\text{log}_{2}(\frac{w_{\text{max}}}{\ell}|E|)\rceil$ bits where $w_{\text{max}}$ is the maximum latency on a link and stores the quotient of the division. For technical reasons, we also have a counter $C_{3}$; its value is always equal to the value of $C_{2}$ plus $1$.
@@
$\exists\pi_{1}.\ \exists\pi_{2}.\\ \mathop{\mathbf{F}\vphantom{a}}\nolimits(\ \pi_{1}.goal\land\pi_{2}.goal\land((\pi_{1}.C_{2}>\pi_{2}.C_{3})\lor(\pi_{1}.C_{2}=\pi_{2}.C_{3}\land\pi_{1}.C_{1}\geq\pi_{2}.C_{1})))\ .$
@@ 5.3 Experimental Setup
Figure 9 shows the CPU time (on y-axis) for all independently sorted instances that where solved by TAPAAL HyperLTL and MCHyper. For each tool, we consider three benchmarks: one with $\ell=8$ and where all latency constants are between 1 to 5 and two more where all constants as well as $\ell$ are multiplied by $15$ respectively $255$. This multiplication creates instances that are equivalent to the basic one, however, the numbers in the model become larger.